OpenStudy (anonymous):

What's 21 quarters out of 186 quarters in simplest form?

OpenStudy (triciaal):

write as fraction 21/186 now find the prime factors that can cancel out 21 = 3 * 7 186 = 3*62 so did you get 7/62?